# Investigating dusty Red Supergiant outflows in Westerlund 1 with 3D Hydrodynamic simulations

### v1

@@ 5 Conclusion
In this work, we computed a 3D hydrodynamic model of a RSG embedded in a hot cluster wind, with values typical for those in Westerlund 1. We summarise our findings below:
@@
The RSG wind material is accelerated as it is heated, and that this acceleration tends towards a steady state over time.
@@
We produce synthetic dust continuum-emission maps for 11 m and 24 m for several viewing angles for comparison with current and future observations of RSG outflows in Westerlund 1 and other clusters.
Our dust maps correspond well to observations of W20, W26, and W237 without any fine-tuning, both in terms of observed fluxes and morphology. Our omission of PAH emission in our dust map modelling suggests that PAHs may be depleted in the ablation flow.
We encourage further observational studies of the fate of RSG winds and dusty clumps in clusters. This will help ascertain the role RSG winds could play in mass loading, as well as their potential to seed future populations of stars as inferred from observations of globular clusters.
Follow-up observations with infrared integral field spectroscopy such as VLT/ERIS-SPIFFIER would enable detailed kinematic studies of the RSG-cluster wind interaction and resulting ablation flow. This would provide insights into the origin of Westerlund 1’s nebular emission, a direct probe of the so-far elusive RSG mass-loss rates and help disentangle the various physical processes in the subsequent mixing.
@@ Acknowledgements.
The simulations presented here were performed on the HPC system Raven at the Max Planck Computing and Data Facility. CJKL gratefully acknowledges support from the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg in the form of an IMPRS PhD fellowship. This publication results from research conducted with the financial support of Taighde Éireann - Research Ireland under Grant number 20/RS-URF-R/3712 (JM). JM is grateful to B. Reville and the Max Planck Institute for Nuclear Physics (MPIK) for welcoming him as a regular visitor during the preparation of this paper. AACS is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). TJH acknowledges funding from a Royal Society Dorothy Hodgkin Fellowship and UKRI guaranteed funding for a Horizon Europe ERC consolidator grant (EP/Y024710/1). This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561. This study made use of the following software packages: Astropy (Astropy Collaboration et al., 2018), Numpy (Harris et al., 2020), matplotlib (Hunter, 2007), yt (Turk et al., 2011), pion (Mackey et al., 2021), pypion (Green & Mackey, 2021), TORUS (Harries et al., 2019).References
- Andersen et al. (2017) Andersen, M., Gennaro, M., Brandner, W., et al. 2017, A&A, 602, A22
- Andrews et al. (2018) Andrews, H., Fenech, D., Prinja, R. K., Clark, J. S., & Hindson, L. 2018, MNRAS, 477, L55
- Arthur (2012) Arthur, S. J. 2012, MNRAS, 421, 1283
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., SipHocz, B. M., et al. 2018, AJ, 156, 123
- Ballone et al. (2018) Ballone, A., Schartmann, M., Burkert, A., et al. 2018, MNRAS, 479, 5288
- Bastian & Lardo (2018) Bastian, N. & Lardo, C. 2018, ARA&A, 56, 83
- Bastian & Strader (2014) Bastian, N. & Strader, J. 2014, MNRAS, 443, 3594
- Beasor et al. (2021) Beasor, E. R., Davies, B., Smith, N., Gehrz, R. D., & Figer, D. F. 2021, ApJ, 912, 16
- Beasor et al. (2020) Beasor, E. R., Davies, B., Smith, N., et al. 2020, MNRAS, 492, 5994
- Candian & Sarre (2015) Candian, A. & Sarre, P. J. 2015, MNRAS, 448, 2960
- Clark et al. (2008) Clark, J. S., Muno, M. P., Negueruela, I., et al. 2008, A&A, 477, 147
- Clark et al. (2005) Clark, J. S., Negueruela, I., Crowther, P. A., & Goodwin, S. P. 2005, A&A, 434, 949
- Clark et al. (2020) Clark, J. S., Ritchie, B. W., & Negueruela, I. 2020, A&A, 635, A187
- Crowther et al. (2006) Crowther, P. A., Hadfield, L. J., Clark, J. S., Negueruela, I., & Vacca, W. D. 2006, MNRAS, 372, 1407
- Davies et al. (2023) Davies, R., Absil, O., Agapito, G., et al. 2023, A&A, 674, A207
- De Colle et al. (2014) De Colle, F., Raga, A. C., Contreras-Torres, F. F., & Toledo-Roy, J. C. 2014, ApJ, 789, L33
- de Mink et al. (2009) de Mink, S. E., Pols, O. R., Langer, N., & Izzard, R. G. 2009, A&A, 507, L1
- Decin (2021) Decin, L. 2021, ARA&A, 59, 337
- Draine (2003) Draine, B. T. 2003, ARA&A, 41, 241
- Fenech et al. (2018) Fenech, D. M., Clark, J. S., Prinja, R. K., et al. 2018, A&A, 617, A137
- Fielding et al. (2020) Fielding, D. B., Ostriker, E. C., Bryan, G. L., & Jermyn, A. S. 2020, ApJ, 894, L24
- Gratton et al. (2012) Gratton, R. G., Carretta, E., & Bragaglia, A. 2012, A&A Rev., 20, 50
- Green & Mackey (2021) Green, S. & Mackey, J. 2021, PyPion: Post-processing code for PION simulation data, Astrophysics Source Code Library, record ascl:2103.026
- Green et al. (2019) Green, S., Mackey, J., Haworth, T. J., Gvaramadze, V. V., & Duffy, P. 2019, A&A, 625, A4
- Gronke & Oh (2020) Gronke, M. & Oh, S. P. 2020, MNRAS, 492, 1970
- Guarcello et al. (2024) Guarcello, M. G., Almendros-Abad, V., Lovell, J. B., et al. 2024, arXiv e-prints, arXiv:2411.13051
- Härer et al. (2023) Härer, L. K., Reville, B., Hinton, J., Mohrmann, L., & Vieu, T. 2023, A&A, 671, A4
- Harries et al. (2019) Harries, T. J., Haworth, T. J., Acreman, D., Ali, A., & Douglas, T. 2019, Astronomy and Computing, 27, 63
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Hartquist et al. (1986) Hartquist, T. W., Dyson, J. E., Pettini, M., & Smith, L. J. 1986, MNRAS, 221, 715
- Haubner et al. (2025) Haubner, K., Sasaki, M., Mitchell, A., et al. 2025, arXiv e-prints, arXiv:2501.12990
- Henney et al. (2009) Henney, W. J., Arthur, S. J., de Colle, F., & Mellema, G. 2009, MNRAS, 398, 157
- H.E.S.S. Collaboration et al. (2022) H.E.S.S. Collaboration, Aharonian, F., Ashkar, H., et al. 2022, A&A, 666, A124
- Hummer (1994) Hummer, D. G. 1994, MNRAS, 268, 109
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Ji et al. (2019) Ji, S., Oh, S. P., & Masterson, P. 2019, MNRAS, 487, 737
- Kavanagh et al. (2011) Kavanagh, P. J., Norci, L., & Meurs, E. J. A. 2011, New A, 16, 461
- Krause et al. (2013) Krause, M., Charbonnel, C., Decressin, T., Meynet, G., & Prantzos, N. 2013, A&A, 552, A121
- Kurucz (1992) Kurucz, R. L. 1992, in IAU Symposium, Vol. 149, The Stellar Populations of Galaxies, ed. B. Barbuy & A. Renzini, 225
- Li et al. (2020) Li, Z., Hopkins, P. F., Squire, J., & Hummels, C. 2020, MNRAS, 492, 1841
- Mackey et al. (2015) Mackey, J., Castro, N., Fossati, L., & Langer, N. 2015, A&A, 582, A24
- Mackey et al. (2021) Mackey, J., Green, S., Moutzouri, M., et al. 2021, MNRAS, 504, 983
- Mackey et al. (2025) Mackey, J., Mathew, A., Ali, A. A., et al. 2025, arXiv e-prints, arXiv:2501.06021
- Mackey et al. (2014) Mackey, J., Mohamed, S., Gvaramadze, V. V., et al. 2014, Nature, 512, 282
- Mathis et al. (1977) Mathis, J. S., Rumpl, W., & Nordsieck, K. H. 1977, ApJ, 217, 425
- Mauron & Josselin (2011) Mauron, N. & Josselin, E. 2011, A&A, 526, A156
- Micelotta et al. (2010) Micelotta, E. R., Jones, A. P., & Tielens, A. G. G. M. 2010, A&A, 510, A36
- Mohamed et al. (2012) Mohamed, S., Mackey, J., & Langer, N. 2012, A&A, 541, A1
- Montargès et al. (2021) Montargès, M., Cannon, E., Lagadec, E., et al. 2021, Nature, 594, 365
- Muno et al. (2006) Muno, M. P., Law, C., Clark, J. S., et al. 2006, ApJ, 650, 203
- Portegies Zwart et al. (2010) Portegies Zwart, S. F., McMillan, S. L. W., & Gieles, M. 2010, ARA&A, 48, 431
- Reiter et al. (2024) Reiter, M., Haworth, T. J., Manara, C. F., et al. 2024, MNRAS, 527, 3220
- Rogers & Pittard (2013) Rogers, H. & Pittard, J. M. 2013, MNRAS, 431, 1337
- Rybicki & Lightman (1979) Rybicki, G. B. & Lightman, A. P. 1979, Radiative processes in astrophysics (John Wiley & Sons, Inc.)
- Siebenmorgen & Krügel (2010) Siebenmorgen, R. & Krügel, E. 2010, A&A, 511, A6
- Silich et al. (2011) Silich, S., Bisnovatyi-Kogan, G., Tenorio-Tagle, G., & Martínez-González, S. 2011, ApJ, 743, 120
- Stevens & Hartwell (2003) Stevens, I. R. & Hartwell, J. M. 2003, MNRAS, 339, 280
- Szécsi et al. (2018) Szécsi, D., Mackey, J., & Langer, N. 2018, A&A, 612, A55
- Tan et al. (2021) Tan, B., Oh, S. P., & Gronke, M. 2021, MNRAS, 502, 3179
- Tappe et al. (2006) Tappe, A., Rho, J., & Reach, W. T. 2006, ApJ, 653, 267
- Turk et al. (2011) Turk, M. J., Smith, B. D., Oishi, J. S., et al. 2011, The Astrophysical Journal Supplement Series, 192, 9
- Vieu & Reville (2023) Vieu, T. & Reville, B. 2023, MNRAS, 519, 136
- Wareing et al. (2007) Wareing, C. J., Zijlstra, A. A., & O’Brien, T. J. 2007, ApJ, 660, L129
- Wesson et al. (2024) Wesson, R., Matsuura, M., Zijlstra, A. A., et al. 2024, MNRAS, 528, 3392
- Westmoquette et al. (2013) Westmoquette, M. S., Dale, J. E., Ercolano, B., & Smith, L. J. 2013, MNRAS, 435, 30
- Westmoquette et al. (2010) Westmoquette, M. S., Slavin, J. D., Smith, L. J., & Gallagher, III, J. S. 2010, MNRAS, 402, 152
- Wiersma et al. (2009) Wiersma, R. P. C., Schaye, J., & Smith, B. D. 2009, MNRAS, 393, 99
- Wright et al. (2014) Wright, N. J., Wesson, R., Drew, J. E., et al. 2014, MNRAS, 437, L1
- Wünsch et al. (2017) Wünsch, R., Palouš, J., Tenorio-Tagle, G., & Ehlerová, S. 2017, ApJ, 835, 60
- Wünsch et al. (2008) Wünsch, R., Tenorio-Tagle, G., Palouš, J., & Silich, S. 2008, ApJ, 683, 683
@@ Appendix A Resolution test
To check the effects of resolution on our results, we ran an otherwise identical model using 1923 grid cells per level of refinement. PION has a minimum injection region radius of 7 cells, making this the lowest achievable resolution without having an artificially large injection region. We compare the time evolution of the RSG wind mass for both models in Fig.10.
For a radius of 0.5 pc, both models have converged to similar values in all panels, and the differences are likely stochastic in origin due to the instabilities in the ablation flow. At 3.0 pc, the instabilities are more poorly resolved, reducing the efficiency of gas mixing. This suggests we do not have sufficient resolution at this radius, making our mass-loading estimate a lower limit. We present additional figures of our low resolution model as Figures 11, 12, 13 and 14.
Fig. 11 shows that the downstream wake contains larger and more contiguous clumps that are more closely confined to the -axis, when comparing this lower-resolution calculation with the higher-resolution simulation in Fig. 1. Despite this qualitative difference between the two resolutions, the phase diagrams (Figs. 12 and 13) nevertheless appear quite similar between the two resolutions. The mass ratio of hot-to-cold gas in Fig. 14(d) is significantly lower than in Fig. 4. This shows that dynamical mixing is not sufficiently well-resolved at this lower numerical resolution, and that higher resolution is required to demonstrate whether or not our results for mass-loading have converged.

### v2

@@ 5 Conclusion
In this work we computed a 3D HD model of a RSG embedded in a hot cluster wind with values typical for those in Westerlund 1. We summarise our findings below:
@@
The RSG wind material is accelerated as it is heated, and this acceleration tends towards a steady state over time.
@@
We produced synthetic dust continuum-emission maps for 11 m and 24 m for several viewing angles for comparison with current and future observations of RSG outflows in Westerlund 1 and other clusters.
Our dust maps closely match observations of W20, W26, and W237 without any fine-tuning, in terms of both observed fluxes and morphology. Our omission of PAH emission in our dust map modelling suggests that PAHs are depleted in the ablation flow due to the effects of winds and radiation and/or a lack of PAH formation in the RSG winds in the first place.
We encourage further observational studies of the fate of RSG winds and dusty clumps in clusters. This will help ascertain the role RSG winds play in mass loading, as well as their potential to seed future populations of stars, as inferred from observations of globular clusters.
Follow-up observations with IR integral field spectroscopy such as with VLT/ERIS-SPIFFIER would enable detailed kinematic studies of the RSG–cluster wind interaction and resulting ablation flow. This would provide insights into the origin of Westerlund 1’s nebular emission, a direct probe of the so-far elusive RSG mass-loss rates, and help disentangle the various physical processes in the subsequent mixing.
@@ Acknowledgements.
We thank the anonymous referee for their constructive feedback which has improved the article. The authors thank Brian Reville for useful discussions on hydrodynamics. The simulations presented here were performed on the HPC system Raven at the Max Planck Computing and Data Facility. CJKL gratefully acknowledges support from the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg in the form of an IMPRS PhD fellowship. This publication results from research conducted with the financial support of Taighde Éireann - Research Ireland under Grant number 20/RS-URF-R/3712 (JM). JM is grateful to B. Reville and the Max Planck Institute for Nuclear Physics (MPIK) for welcoming him as a regular visitor during the preparation of this paper. AACS is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group – Project-ID 445674056 (SA4064/1-1, PI Sander). TJH acknowledges funding from a Royal Society Dorothy Hodgkin Fellowship and UKRI guaranteed funding for a Horizon Europe ERC consolidator grant (EP/Y024710/1). This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561. This study made use of the following software packages: Astropy (Astropy Collaboration et al., 2018), Numpy (Harris et al., 2020), matplotlib (Hunter, 2007), yt (Turk et al., 2011), pion (Mackey et al., 2021), pypion (Green & Mackey, 2021), TORUS (Harries et al., 2019).Contributions. CJKL and JM conceived the project. CJKL performed and analysed the simulations and led the writing of the article. JM computed the TORUS dust maps, assisted with performing the simulations and interpreting the results. TJH created the JWST figures, assisted with TORUS and interpreting the results. AACS provided feedback on the manuscript and theoretical support for the article.
@@ Appendix A Resolution test
To check the effects of resolution on our results, we ran an otherwise identical model using 1923 grid cells per level of refinement. PION has a minimum injection region radius of 7 cells, making this the lowest achievable resolution without having an artificially large injection region. We compare the time evolution of the RSG wind mass for the two models in Fig.10.
For a radius of 0.5 pc, both models have converged to similar values in all panels, and the differences likely originate from instabilities in the ablation flow. At 3.0 pc, the instabilities are more poorly resolved, reducing the efficiency of gas mixing. This suggests we do not have sufficient resolution at this radius, making our mass-loading estimate a lower limit. We present additional figures of our low resolution model as Figs. 11, 12, 13, and 14.
Figure 11 shows that the downstream wake contains larger and more contiguous clumps that are more closely confined to the -axis, when comparing this lower-resolution calculation with the higher-resolution simulation in Fig. 1. Despite this qualitative difference between the two resolutions, the phase diagrams (Figs. 12 and 13) nevertheless appear quite similar between the two resolutions. The mass ratio of hot-to-cold gas in Fig. 14(d) is significantly lower than in Fig. 4. This shows that dynamical mixing is not sufficiently well resolved at this lower numerical resolution, and that higher resolution is required to demonstrate whether or not our results for mass loading have converged.